OpenStudy (alprincenofl):

Find f(1),f(2),f(3) and f(4), if f(n) is defined recursively by f(0)=1 and for n=0,1,2,3,… A) f(n+1)=f(n)+2 B) f(n+1)=2^f(n) .

OpenStudy (alprincenofl):

http://prntscr.com/57tzjm

OpenStudy (anonymous):

this is two separate questions right?

OpenStudy (anonymous):

\[f(0)=1\\ f(n+)=f(n)+2\]

OpenStudy (anonymous):

so \[f(1)=f(0)+2=0+2=2\] and \[f(2)=f(1)=2+2=4\] and so on